OpenStudy (perl):

Algebra problem , prove the following is an identity: x* ( 1 + x)^7/7 - (1 + x )^8/(7*8) = (1+x)^8/8 - (1+x)^7/7 You can start from the left side or right side. (the problem arose when integrating x (1+x)^6 , i solved it two different ways using u sub and integration by parts).

OpenStudy (perl):

i made various attempts and got stuck

OpenStudy (perl):

Attempt: (1+x)^8/8 - (1+x)^7/7 = ( 1 + x)^7 [ ( 1 + x)/8 - 1/7 ] = ( 1 + x)^7 [ ( 7 + 7x - 8 ) / (8*7) ] = (1+x)^7 [ ( 7x - 1 ) / (8*7) ] =. . . = x* ( 1 + x)^7/7 - (1 + x )^8/(7*8)

OpenStudy (michele_laino):

left side: \[=\frac{ 8x(1+x)^{7}-(1+x ^{8}) }{ 56 }=\frac{ (1+x)^{7}(7x-1) }{ 56 }\] right side: \[=\frac{ 7(1+x)^{8}-8(1+x)^{7} }{ 56 }=\frac{ (7x-1)(1+x)^{7} }{ 56 }\]

OpenStudy (perl):

that works :)

OpenStudy (perl):

you formed a bridge between right hand side and left hand side

OpenStudy (michele_laino):

yes! I solved your identity like a trigonometric identity!

OpenStudy (perl):

so it is easier to show that both sides simplify to the same expression, and then go from there

OpenStudy (michele_laino):

yes!

OpenStudy (perl):

didn't think of that :)

OpenStudy (michele_laino):

thanks!

OpenStudy (anonymous):

I MEANT ON MESSAGE.....

OpenStudy (anonymous):

plz message them to me

OpenStudy (anonymous):

Subtract the two sides and show that it is zer0 \[-\frac{1}{8} (x+1)^8-\frac{(x+1)^8}{7\times 8}+\frac{1}{7} (x+1)^7+\frac{1}{7} (x+1)^7 x=\\ -\frac{1}{7} (x+1)^8+\frac{1}{7} (x+1)^7+\frac{1}{7} x (x+1)^7=\\ \frac{1}{7} (x+1)^7 \left(\frac{x}{7}-\frac{x}{7}+\frac{1}{7}-\frac{1}{7}\right)=0\]

OpenStudy (perl):

yes that works too :) I thought there was a simple way to factor the right hand side of the original equation so that it would turn into the left hand side

OpenStudy (perl):

but this is an acceptable way to prove an identity, as well

OpenStudy (anonymous):

\[\text {One Side} = \frac {1} {7} x (x + 1)^7 - \frac {1} {56} (x + 1)^8 = \\\frac {8} {57} x (x + 1)^7 - \frac {1} {56} (x + 1)^8 = \\ \frac {1} {56} (x + 1)^7 (8 x - x - 1) = \frac {1} {56} (x + 1)^7 (7 x - 1)\]

OpenStudy (anonymous):

\[\text {Other Side} = \frac {1} {8} (x + 1)^8 - \frac {1} {7} (x + 1)^7 = \\\frac {7} {56} (x + 1)^8 - \frac {8} {56} (x + 1)^7 =\\ \frac {1} {56} (x + 1)^7 (7 (x + 1) - 8) = \frac {1} {56} (x + 1)^7 (7 x - 1)\]

OpenStudy (anonymous):

@perl